The following data were collected during a short race between

two friends.

Velocity (m/s) 0 0.5 1 1.5 2 2 4 6 2 0

Time (s) 0 2 4 6 8 10 12 14 16 18

a) Describe the different sections of the graph.

b) Determine the acceleration over the first eight seconds.

c) Determine the maximum acceleration.

d) Using the graph calculate the displacement:

i) over the first eight seconds

ii) the total race.

e) Find the maximum velocity reached by the runner.


1
Expert's answer
2021-12-06T09:46:55-0500


a) We have accelerated motion (0 to 8 s), uniform motion (8 to 10 s), more accelerated (10 to 14), decelerated (14 to 16), a bit less decelerated (16 to 18).

b) The acceleration is


"a=\frac vt=0.25\text{ m/s}^2."

c) The maximum:


"a=\frac vt=2\text{ m/s}^2."

d) Displacement:


"d(8)=8·2/2=8\text{ m}.\\
d_{18}=38\text{ m}."

e) Max. velocity:


"v_\text{max}=v(14)=6\text{ m/s}."